A weather balloon is launched with a volume of 32 L at SATP.
Fifteen minutes later, the balloon has a volume of 45 L, a temperature of -24.5 °C.
What is the final pressure of the balloon?
1
Expert's answer
2020-04-06T11:51:07-0400

According to the ideal gas law:


"\frac{P_{1}V_{1}}{T_{1}}=\frac{P_{2}V_{2}}{T_{2}},"


or


"P_{2}=\frac{P_{1}V_{1}T_{2}}{T_{1}V_{2}}=\frac{101.3kPa*32L*248.5K}{45L*273K}=65.6kPa(0.65atm)."